OpenStudy (anonymous):

Find the domain and range of the function. h(x)= - (square root of x+3)

OpenStudy (anonymous):

f(x) = √(x-3) f(x+h) = √(x+h-3) f(x+h) - f(x) = √(x+h-3) - √(x-3) [ f(x+h) - f(x) ] /h = [√(x+h-3) - √(x-3)]/h multiply top and bottom with [√(x+h-3) + √(x-3)] [ f(x+h) - f(x) ] /h = [(x+h-3) - (x-3)]/h[√(x+h-3) + √(x-3)] = [ h /h[√(x+h-3) + √(x-3)] = [ 1 /[√(x+h-3) + √(x-3)] when h--> 0 derivative of √(x-3 = 1 / [√(x-3) + √(x-3)] = 1/(2√(x-3))

OpenStudy (anonymous):

so thats the domain and the range?